OpenStudy (moongazer):

teach me on how to get the value of this 3 variables I have slight knowledge about this but i can't get the value with these systems of linear equations help me please Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (moongazer):

2x+3y-z=4 3x-4y+4z=11 5x-2y+z=5

OpenStudy (anonymous):

let's cancel a variable

OpenStudy (anonymous):

Setup the augmented matrix: \[\left[\begin{matrix}2 & 3 & -1 & 4 \\ 3 & -4 & 4 & 11 \\ 5 & -2 & 1 & 5\end{matrix}\right]\] Perform row reduction (gaussian elimination): http://en.wikipedia.org/wiki/Gaussian_elimination Add -3/2 times row 1 to row 2. Add -5/2 times row 1 to row 3. Add -19/17 times row 2 to row 3. Scale row 3 by -17/45. Add -11/2 times row 3 to row 2. Add 1 times row 3 to row 1. Scale row 2 by -2/17. Add -3 times row 2 to row 1. Scale row 1 by 1/2. Result: \[\left[\begin{matrix}1 & 0 & 0 & 1 \\ 0 & 1 & 0 & 2 \\ 0 & 0 & 1 & 4\end{matrix}\right]\] x = 1 y = 2 z = 4

OpenStudy (moongazer):

could you use the long method? because i think matrices are confusing

OpenStudy (anonymous):

Let's cancel z

OpenStudy (moongazer):

yup

OpenStudy (moongazer):

then?

OpenStudy (anonymous):

Add -3/2 times equation 1 to equation 2. Add -5/2 times equation 1 to equation 3. Add -19/17 times equation 2 to equation 3. Multiply both sides of equation 3 by -17/45. Add -11/2 times equation 3 to equation 2. Add 1 times equation 3 to equation 1. Multiply both sides of equation 2 by -2/17. Add -3 times equation 2 to equation 1. Multiply both sides of equation 1 by 1/2.

OpenStudy (callisto):

2x+3y-z=4 z = 2x+3y-4 --(1) 3x-4y+4z=11 => 3x-4y+4z=11 --(2) 5x-2y+z=5 5x-2y+z=5 --(3) Sub(1) into (2) and (3) respectively 3x-4y+4(2x+3y-4)=11 =>11x+8y -16 =11 5x-2y+(2x+3y-4)=5 7x+y-4 =5 11x+8y =27 –-(4) y =9-7x –-(5) Sub. (5) into (4) 11x+8(9-7x) =27 => -45x=-45 => x=1 Sub x=1 into (5) y =9-7(1) = 2 Sub x=1 and y=2 into (1) z = 2(1)+3(2)-4 =4 so x=1, y=2 , z=4

OpenStudy (anonymous):

2x+3y-z=4 3x-4y+4z=11 5x-2y+z=5 Multiply 1st equation by 4, third by -4, and add canceling z -3x+6y=9 can be simplified into -x+2y=3 2y=x+3 Plug that back into equation 5x-2y+z=5 5x-(x+3)+z=5 4x-3+z=5 4x+z=8 3x-4y+4z=11 3x-2(x+3)+z=11 3x-2x+6+z=11 x+z=5 4x+z=8 x+z=5 Multiply second equ by -1 and add two equation 3x=3 x=1 x+z=5 1+z=5 z=4 let's find y using this equation from above 2y=x+3 2y=1+3 2y=4 y=2

OpenStudy (moongazer):

i'll try to work this in a paper

OpenStudy (moongazer):

confusing

OpenStudy (moongazer):

step by step please

OpenStudy (anonymous):

Try my way, it includes pretty much every steps

OpenStudy (moongazer):

it confused me

OpenStudy (callisto):

lol, it takes much more time.If you know how to solve system of equations in 2 unknowns. then it's easier. In this case just make one of the variable in the equation as the subject and sub into other equations. then you'll get 2 equations in 2 unknowns then solve them

OpenStudy (moongazer):

i know how to solve 2 unknowns

OpenStudy (anonymous):

For anything larger than a 3x3 you should really use row reduction.

OpenStudy (moongazer):

my teacher didn't taught us that

OpenStudy (callisto):

if you know how to solve 2unknown in 2 equations then, just make one of the variables into the subject. like i have made z as teh subject and sub into teh other equations then i get 2equations in 2unknowns In this case you can solve it. first solve x and y and sub the value of x and y to solve z :)

OpenStudy (callisto):

sorry teh = the, typing mistakes

OpenStudy (moongazer):

this is what i made: 2x+3y-z=4 (1) 3x-4y+4z=11 (2) 5x-2y+z=5 (3) I multiplied eq. 1 by 4 (2x+3y-z=4)4 then added it to eq.2 8x+12y-4z=16 (1) 3x-4y+4z=11 (2) =11x+8y=27 (4) then i multiplied eq.3 to -4 and added to eq.2 (5x-2y +z=5)4 then added it to eq.2 3x-4y+4z=11 eq.2 -20x+8y-4z=-20 eq. 3 =-17+4y=9 (5) then i multiplied (5) to -2 34x-8y=-18 (5) then i added it to (4) 11x+8y=27 34x-8y=-18 = 45x=9 x=9/45 i know my answer is wrong, but where did i go wrong?

OpenStudy (callisto):

=-17+4y=9 (5) it should be -9 instead of 9

OpenStudy (callisto):

change the value and keep your steps, then ''then i multiplied (5) to -2 34x-8y=18 (5) then i added it to (4) 11x+8y=27 34x-8y=18 45x=45 x=1 ''

OpenStudy (moongazer):

wow that is a simple mistake that made me a failing quiz. Signs always get me in trouble, even in my exam I got a mistake because of a wrong sign. I will be careful know with my signs. ^_^

OpenStudy (moongazer):

Thanks!